# The Relative Contributions of Accretion Disk versus Jet to the Optical and Mid-infrared Variability of Seyfert Galaxies

### v1

@@ Abstract
We performed a comprehensive analysis of flux and color variability in a redshift-matched sample of Seyfert galaxies, comprising 23 gamma-ray-detected narrow-line Seyfert 1 galaxies (gNLS1s), 190 non-gamma-ray-detected narrow-line Seyfert 1 galaxies (ngNLS1s), and 10 gamma-ray-detected broad-line Seyfert 1 galaxies (gBLS1s). Utilizing multi-band light curves from the Zwicky Transient Facility (ZTF) in g, r, and i bands, along with mid-infrared (MIR) observations in W1 and W2 bands from the Wide-field Infrared Survey Explorer (WISE), we observed that gBLS1s exhibit more significant variability than gNLS1s, while ngNLS1s display minimal variability across both optical and MIR wavelengths. The pronounced variability in gBLS1s may be attributed to a more closely aligned jet relative to the observer’s line of sight or their comparatively lower accretion rates. In contrast, the subdued variability in ngNLS1s suggests that their flux changes are primarily driven by accretion disk instabilities. A significant correlation between optical and MIR variability amplitudes found here supports the reprocessing scenario, wherein variations in the accretion disk emission are re-emitted by surrounding dust. Furthermore, our long-term color variability analysis revealed a stronger redder-when-brighter (RWB) trend in approximately 35%, 61%, and 80% of gNLS1s, ngNLS1s, and gBLS1s, respectively, in optical wavelength, strengthens the reprocessing scenario with the observed trend of stronger RWB for approximately 68%, 96%, and 30%, respectively, in MIR wavelength. The prevalent RWB trend observed in both optical and MIR wavelengths from the current sample of Seyfert galaxies on the longer time scales is likely associated with accretion disk instabilities.
@@
Among the defining characteristics of AGNs, flux variability across the electromagnetic spectrum stands out as one of the most significant traits (Wagner & Witzel, 1995; Ulrich et al., 1997). Such variability, observed on timescales ranging from hours to years across wavelengths from radio to -ray, has been extensively utilized to investigate emitting regions, the intricate nature of the central engine, and accretion processes, phenomena (see, Urry & Padovani, 1995; Wagner & Witzel, 1995; Ulrich et al., 1997; Czerny et al., 2008; Rakshit et al., 2015). Among these, optical variability has been a main focus of study since the discovery of AGNs, as it provides crucial insights into the physics of accretion disks and jets (Wagner & Witzel, 1995; Ghosh et al., 2000; Villata et al., 2004; Gupta et al., 2008; Rani et al., 2010; Gu & Ai, 2011; Ikejiri et al., 2011; Bonning et al., 2012; Mao & Zhang, 2016). In addition to optical studies, the investigation of infrared (IR) variability has proven equally important. Infrared echoes, which arise in response to variations in optical continuum, provide vital information about the morphology of the dust torus and the accretion mechanisms that power the AGNs (Maune et al., 2013). Moreover, IR wavelength presents unique advantages over optical wavelength, such as (1) the IR continuum being unaffected by the presence of strong emission lines in low-redshift AGNs, and (2) reduced sensitivity to dust extinction. Together, multiwavelength variability studies across the optical and IR bands provide a comprehensive understanding of the central regions and the processes shaping AGN emission. In particular, changes in the brightness of AGNs are often accompanied by spectral variations, which can be explored through color-magnitude or spectral index–magnitude correlations (e.g., see Ciprini et al., 2003). Examining spectral variability is particularly important, as it can help to pinpoint the origin of brightness fluctuations, even when the analysis is confined to the optical and IR bands (Vagnetti et al., 2003).
@@
Optical variability studies on large samples of luminous AGNs in optical wavelength have resulted in various interesting relationships between the amplitude and timescale of variability with physical parameters of AGN such as black hole mass, Fe ii strength, accretion rate, bolometric luminosity, redshift, etc. (e.g., di Clemente et al., 1996; Vanden Berk et al., 2004; Kelly et al., 2009; MacLeod et al., 2010; Meusinger et al., 2011; Rakshit & Stalin, 2017; Li et al., 2018), providing new insights into the accretion process in AGNs. However, it has been limited to a few sources only for the low luminous part of AGNs such as narrow-line Seyfert 1 galaxies (NLS1s), characterized based upon the flux ratio of forbidden line with Balmer line (H) 3, and narrow width of Balmer emission line with full width at half maximum (FWHM) 2000 km s-1 (Osterbrock & Pogge, 1985; Goodrich et al., 1989). For instance, a few studies on radio-loud and radio-quiet NLS1s have been performed on the intra-night timescale (Paliya et al., 2013; Kshama et al., 2017; Ojha, 2022; Ojha et al., 2022, 2024) but very limited studies are available on longer time scales from days to year (Young et al., 1999; Miller et al., 2000; Klimek et al., 2004; Doroshenko et al., 2006; Maune et al., 2013). Similarly, scarce studies are available for NLS1s in the IR band (Jiang et al., 2012). On the other hand, NLS1s have a higher accretion rate, strong Fe ii emission, stronger X-ray variability, steeper X-ray spectra, lower black hole mass, and lower optical variability compared to their broad-line counterparts, namely broad-line Seyfert 1 galaxies (BLS1s; Leighly, 1999; Véron-Cetty et al., 2001; Grupe & Mathur, 2004; Zhou et al., 2006; Xu et al., 2012; Ojha et al., 2020). Despite the intrinsic differences between NLS1s and BLS1s, gamma-ray-detected Seyfert galaxies, including both NLS1s and BLS1s, are of particular interest because they exhibit optical variability on short timescales, similar to that observed in blazars (Ojha et al., 2022). This suggests a significant contribution from relativistic jets in these sources, raising the question whether the observed optical and MIR variability arises primarily from jet activity or from thermal processes in the accretion disk.
@@ 2 Sample selection
For this study, we used the most recent catalog of NLS1 galaxies, comprising 22,656 sources (Paliya et al., 2024). To construct a sample of ngNLS1s, we first compiled all known gNLS1s identified by the Fermi Large Area Telescope (LAT) from the literature, including the 21 sources listed in Table 1 of Paliya et al. (2024). This compilation yielded a total of 25 gNLS1s. Applying a redshift cut of to ensure adequate signal-to-noise ratios (S/N) for variability analysis using data from the ZTF and WISE reduced the gNLS1 sample to 23 sources. In parallel, a control sample of 45 gBLS1s was selected from Paliya et al. (2024), also subject to the same redshift constraint of to maintain consistency in data quality across samples. Thus, finally, we identified 22,635 ngNLS1s, 23 gNLS1s, and 45 gBLS1s. Next, we prepared redshift-matched samples to enable a fair comparison. The resulting matched samples consist of 23 gNLS1s, 190 ngNLS1s, and 10 gBLS1s whose cumulative probability distribution for rest-frame redshifts is presented in Fig. 1 . Finally, a Kolmogorov–Smirnov (K-S) test performed on the redshift-matched samples yielded a null hypothesis probability () of approximately 21% for gNLS1s versus ngNLS1s and 28% for gNLS1s versus gBLS1s.
Out of different subsamples of Seyfert galaxies used in the current work, gNLS1s and gBLS1s can exhibit hour-like time-scale variability, which can create an obstacle in studying the true spectral behavior of these sources in the absence of simultaneous multi-band data, both in optical and MIR wavebands. Therefore, for the color variability study, we selected sources that have quasi-simultaneous observations within 30 minutes in different bands with at least five data points. To make it consistent, we have applied the same criterion of quasi-simultaneous observations to the sample of ngNLS1s. This has resulted in a sample of 17 gNLS1s, 128 ngNLS1s, and 10 gBLS1s for the g-r color study and 10 gNLS1s, 74 ngNLS1s, and 6 gBLS1s for the r-i color study in the optical wavelength. Similarly, samples of 22 gNLS1s, 176 ngNLS1s, and 10 gBLS1s resulted in for the W1-W2 color study in the mid-infrared wavelength.
@@ 2.1 Optical photometric data from ZTF
For the redshift-matched samples of 23 gNLS1s, 190 ngNLS1s, and 10 gBLS1s, we retrieved ZTF photometric data for the g, r, and i band in the 22nd ZTF public data release (see Masci et al., 2019) within a search radius of an arcsec to the target position from the NASA/IPAC IR Science Archive (IRSA111https://irsa.ipac.caltech.edu/Missions/ztf.html) using the application program interface (API). To eliminate any suspicious measurements that could mimic spurious variability, such as those arising from the independent calibration of ZTF light curves across different fields and CCD quadrants for the same filter (e.g., see van Roestel et al., 2021), we select only the light curve for a given filter corresponding to the observation ID with the maximum number of data points. Furthermore, to ensure the quality of the photometric data, we selected data points with catflags = 0, as recommended in the ZTF Science Data System Explanatory Supplement222https://irsa.ipac.caltech.edu/data/ZTF/docs/ztf explanatory supplement.pdf and the public data release notes333https://irsa.ipac.caltech.edu/data/ZTF/docs/releases/dr20/ztf release notes dr20.pdf, removed outliers by applying a 3 clipping on the entire data set and also excluded poor quality data points with uncertainties greater than 10% in magnitude. The precautions and filtering above enable us to determine small-magnitude variations in the generated g, r, and i band light curves. This has resulted in samples of 20 gNLS1s, 186 ngNLS1s, and 10 gBLS1s for the ZTF r-band. The total number of sources with available data in ZTF g, r, and i bands for different classes of Seyfert galaxies is tabulated in Table 1. The obtained lightcurves of Seyfert galaxies have magnitude data points at a cadence of as high as 1 day for sources in their field with a maximum time baseline of 2000 days. Representative light curves of the r-band for one of the gNLS1-J032441.28341045.1 (left), ngNLS1-J221541.64301435.7 (middle) and gBLS1- J095649.88251516.1 (right) from the current sample are shown in Fig. 2.
@@ 2.2 Mid-infrared data from WISE
For the redshift-matched samples of 23 gNLS1s, 190 ngNLS1s, and 10 gBLS1s, we retrieved MIR data from the MEP database and the NEOWISE-R single exposure (L1b) source database from the NASA/IPAC IR Science Archive444https://irsa.ipac.caltech.edu/Missions/wise.html, giving us data from 2010 to 2024 with a gap of about 3 years between 2011-2014 due to the depletion of the solid hydrogen cryogen used to cool the W3 and W4 instruments. Next, we limited our MIR study on the current sample to the W1 and W2 bands only due to the availability of the data for the W3 and W4 bands between 2010 and 2011, with many photometric points marked as ‘null’. Furthermore, both databases were screened to exclude bad photometric measurements, adopting the criteria (i) the reduced value of for the single exposure profile fit must be below 5 for both the W1 and W2 bands (that is, w1rchi2 5 and w2rchi2 5). (ii) The number of points spread function (PSF) components used in the profile fit for a source should be fewer than 3 (nb 3). (iii) Single-exposure images must have the highest quality (qifact = 1), be free from known artifacts (ccflags = ‘0000’), and not involve active d-eblending (na = 0). (iv) The number of photometric data points available in the W1 and W2 bands to be 5. (v) Also removed outliers by applying a 3 clipping on the whole data set.
@@ 3 Methodology for analysis
Before proceeding further with the analysis, we first corrected the observed data of each source for the foreground galactic interstellar reddening and absorption, using the extinction correction (AV) values provided by NASA’s NED, following the method of (Schlafly & Finkbeiner, 2011). The extinction-corrected r-band light curves for a gNLS1, ngNLS1, and gBLS1 from the current sample are displayed in Fig. 2.
@@ 3.1 Optical flux variability
To confirm the presence/absence of variability in the final samples of gNLS1s, ngNLS1s, and gBLS1s (see Table 1) using the ZTF g, r, and i band light curves, we have used the standard -test (see, Goyal et al., 2012). A detailed explanation for this test is demonstrated in our previous papers (Ojha et al., 2022, 2024, and references therein).
In brief, following Goyal et al. (2012), -test can be written as
@@
where, is the mean square error and given as follows:
is the sample variance for the flux measurements, . The associated error with can be computed following the given equation
For each gNLS1, ngNLS1, and gBLS1 galaxies, we have tabulated the duration of the light curve, number of data points in the light curve, variability status based on the -test, , and in Table 2.
@@ 3.1.1 Duty cycle of variability
AGNs do not necessarily exhibit flux variations during each epoch of their monitoring. Hence, it is more suitable to evaluate the duty cycle (DC) as the ratio of the time intervals when variability is observed to the total observing time rather than simply considering the fraction of variable objects. Following the definition proposed by Romero et al. (1999) and further utilized by Stalin et al. (2004) for intra-night variability, we computed the DC for the current sample of Seyfert galaxies using the following expression.
@@ 3.2 Mid-infrared flux variability
where is the observed magnitude at the point, and is the weighted average. The can be expressed as:
where is calculated including systematic error () with the individual error () as follows:

### v2

@@ Abstract
We performed a comprehensive analysis of flux and color variability in a redshift-matched sample of Seyfert galaxies, comprising 23 gamma-ray-detected narrow-line Seyfert 1 galaxies (gNLS1s), 190 non-gamma-ray-detected narrow-line Seyfert 1 galaxies (ngNLS1s), and 10 gamma-ray-detected broad-line Seyfert 1 galaxies (gBLS1s). Utilizing multi-band light curves from the Zwicky Transient Facility (ZTF) in g, r, and i bands, along with mid-infrared (MIR) observations in W1 and W2 bands from the Wide-Field Infrared Survey Explorer (WISE), we observed that gBLS1s exhibit more significant variability than gNLS1s, while ngNLS1s display minimal variability across both optical and MIR wavelengths. The pronounced variability in gBLS1s may be attributed to a more closely aligned jet relative to the observer’s line of sight or their comparatively lower accretion rates. In contrast, the subdued variability in ngNLS1s suggests that their flux changes are primarily driven by accretion disk instabilities. A strong correlation between optical and MIR variability amplitudes across different time scales supports the reprocessing scenario, where accretion disk emission variations are re-emitted by surrounding dust. Furthermore, our long-term color variability analysis revealed both stronger bluer-when-brighter (BWB) and redder-when-brighter (RWB) trends from the current sample, but a stronger RWB in approximately 50%, 49%, and 50% of gNLS1s, ngNLS1s, and gBLS1s, respectively, in the longer side of the optical wavelength, and 55%, 28%, and 30% in the MIR wavelength, strengthen the reprocessing scenario. The prevalent RWB trend observed in both optical and MIR wavelengths from the current sample on the longer time scales is likely associated with accretion disk instabilities.
@@
Among the defining characteristics of AGNs, flux variability across the electromagnetic spectrum stands out as one of the most significant traits (Wagner & Witzel, 1995; Ulrich et al., 1997). Such variability, observed on timescales ranging from hours to years across wavelengths from radio to -ray, has been extensively utilized to investigate emitting regions, the intricate nature of the central engine, and accretion processes, phenomena (see, Urry & Padovani, 1995; Wagner & Witzel, 1995; Ulrich et al., 1997; Czerny et al., 2008; Rakshit et al., 2015). Among these, optical variability has been a main focus of study since the discovery of AGNs, as it provides crucial insights into the physics of accretion disks and jets (Wagner & Witzel, 1995; Ghosh et al., 2000; Villata et al., 2004; Gupta et al., 2008; Rani et al., 2010; Gu & Ai, 2011; Ikejiri et al., 2011; Bonning et al., 2012; Mao & Zhang, 2016). In addition to optical studies, the investigation of infrared (IR) variability has proven equally important. Infrared echoes, which arise in response to variations in optical continuum, provide vital information about the morphology of the dust torus and the accretion mechanisms that power the AGNs (Maune et al., 2013). Moreover, IR wavelengths present unique advantages over optical wavelengths, such as (1) the IR continuum being unaffected by the presence of strong emission lines in low-redshift AGNs, and (2) reduced sensitivity to dust extinction. Together, multiwavelength variability studies across the optical and IR bands provide a comprehensive understanding of the central regions and the processes shaping AGN emission. In particular, changes in the brightness of AGNs are often accompanied by spectral variations, which can be explored through color-magnitude or spectral index–magnitude correlations (e.g., see Ciprini et al., 2003). Examining spectral variability is particularly important, as it can help to pinpoint the origin of brightness fluctuations, even when the analysis is confined to the optical and IR bands (Vagnetti et al., 2003).
@@
Optical variability studies on large samples of luminous AGNs in optical wavelength have resulted in various interesting relationships between the amplitude and timescale of variability with physical parameters of AGN such as black hole mass, Fe ii strength, accretion rate, bolometric luminosity, redshift, etc. (e.g., di Clemente et al., 1996; Vanden Berk et al., 2004; Kelly et al., 2009; MacLeod et al., 2010; Meusinger et al., 2011; Rakshit & Stalin, 2017; Li et al., 2018), providing new insights into the accretion process in AGNs. However, it has been limited to a few sources only for the low luminous part of AGNs such as narrow-line Seyfert 1 galaxies (NLS1s), characterized based upon the flux ratio of forbidden line with Balmer line (H) 3, and narrow width of Balmer emission line with full width at half maximum (FWHM) 2000 km s-1 (Osterbrock & Pogge, 1985; Goodrich et al., 1989). For instance, a few studies on radio-loud and radio-quiet NLS1s have been performed on the intra-night timescale (Paliya et al., 2013; Kshama et al., 2017; Ojha, 2022; Ojha et al., 2022; Ojha et al., 2024) but very limited studies are available on longer time scales from days to year (Young et al., 1999; Miller et al., 2000; Klimek et al., 2004; Doroshenko et al., 2006; Maune et al., 2013). Similarly, scarce studies are available for NLS1s in the IR band (Jiang et al., 2012). On the other hand, NLS1s have a higher accretion rate, strong Fe ii emission, stronger X-ray variability, steeper X-ray spectra, lower black hole mass, and lower optical variability compared to their broad-line counterparts, namely broad-line Seyfert 1 galaxies (BLS1s; Leighly, 1999; Véron-Cetty et al., 2001; Grupe & Mathur, 2004; Zhou et al., 2006; Xu et al., 2012; Ojha et al., 2020). Despite the intrinsic differences between NLS1s and BLS1s, gamma-ray-detected Seyfert galaxies, including both NLS1s and BLS1s, are of particular interest because they exhibit optical variability on short timescales, similar to that observed in blazars (Ojha et al., 2022). This suggests a significant contribution from relativistic jets in these sources, raising the question whether the observed optical and MIR variability arises primarily from jet activity or from thermal processes in the accretion disk.
@@ 2 Sample selection
For this study, we used the most recent catalog of NLS1 galaxies, comprising 22,656 sources (Paliya et al., 2024). To construct a sample of ngNLS1s, we first compile all known gNLS1s identified by the Fermi Large Area Telescope (LAT) from the literature, including the 21 sources listed in Table 1 of Paliya et al. (2024). This compilation yielded a total of 25 gNLS1s. Applying a redshift cut of to ensure adequate signal-to-noise ratios (S/N) for the variability analysis using data from ZTF and WISE reduced the gNLS1 sample to 23 sources. In parallel, a control sample of 45 gBLS1s was selected from Paliya et al. (2024), also subject to the same redshift constraint of to maintain consistency in data quality across samples. Thus, finally, we identified 22,635 ngNLS1s, 23 gNLS1s, and 45 gBLS1s. Next, we prepared redshift-matched samples to enable a fair comparison. The resulting matched samples consist of 23 gNLS1s, 190 ngNLS1s, and 10 gBLS1s whose cumulative probability distribution for rest-frame redshifts is presented in Fig. 1. Finally, a Kolmogorov–Smirnov (K-S) test performed on the redshift-matched samples yielded a null hypothesis probability () of approximately 21% for gNLS1s versus ngNLS1s and 28% for gNLS1s versus gBLS1s.
Out of different subsamples of Seyfert galaxies used in the current work, gNLS1s and gBLS1s can exhibit hour-like time-scale variability, which can create an obstacle in studying the true spectral behavior of these sources in the absence of simultaneous multi-band data, both in optical and MIR wavebands. Therefore, for the color variability study, we selected sources that have quasi-simultaneous observations within 30 minutes in different bands with at least five data points. To make it consistent, we have applied the same criterion of quasi-simultaneous observations to the sample of ngNLS1s. This has resulted in a sample of 17 gNLS1s, 128 ngNLS1s, and 10 gBLS1s for the g-r color study and 10 gNLS1s, 74 ngNLS1s, and 6 gBLS1s for the r-i color study in the optical wavelength. Similarly, samples of 22 gNLS1s, 176 ngNLS1s, and 10 gBLS1s were obtained for the W1-W2 color study at the mid-infrared wavelength.
@@ 2.1 Optical photometric data from ZTF
For the redshift-matched samples of 23 gNLS1s, 190 ngNLS1s, and 10 gBLS1s, we retrieved ZTF photometric data for the g, r, and i band in the 22nd ZTF public data release (see Masci et al., 2019) within a search radius of one arcsec to the target position from NASA/IPAC IR Science Archive (IRSA111https://irsa.ipac.caltech.edu/Missions/ztf.html) using the application program interface (API). To eliminate any suspicious measurements that could mimic spurious variability, such as those arising from the independent calibration of ZTF light curves across different fields and CCD quadrants for the same filter (e.g., see van Roestel et al., 2021), we select only the light curve for a given filter corresponding to the observation ID with the maximum number of data points. Furthermore, to ensure the quality of the photometric data, we selected data points with catflags = 0, as recommended in the ZTF Science Data System Explanatory Supplement222https://irsa.ipac.caltech.edu/data/ZTF/docs/ztf explanatory supplement.pdf and the public data release notes333https://irsa.ipac.caltech.edu/data/ZTF/docs/releases/dr20/ztf release notes dr20.pdf, removed outliers by applying a 3 clipping on the entire data set and also excluded poor quality data points with uncertainties greater than 10% in magnitude. The precautions and filtering above enable us to determine small-magnitude variations in the generated g, r, and i band light curves. This has resulted in samples of 20 gNLS1s, 186 ngNLS1s, and 10 gBLS1s for the ZTF r band. The total number of sources with available data in the ZTF g, r, and i bands for different classes of Seyfert galaxies is tabulated in Table 1. The obtained lightcurves of Seyfert galaxies have magnitude data points at a cadence of as high as 1 day for sources in their field with a maximum time baseline of 2000 days. Representative light curves in the r band for one of the gNLS1-J032441.28341045.1 (left), ngNLS1-J221541.64301435.7 (middle) and gBLS1- J095649.88251516.1 (right) from the current sample are shown in Fig. 2.
@@ 2.2 Mid-infrared data from WISE
For the redshift-matched samples of 23 gNLS1s, 190 ngNLS1s, and 10 gBLS1s, we retrieved MIR data from the MEP database and the NEOWISE-R single exposure (L1b) source database from the NASA/IPAC IR Science Archive444https://irsa.ipac.caltech.edu/Missions/wise.html, giving us data from 2010 to 2024 with a gap of approximately 3 years between 2011 and 2014 due to the depletion of the solid hydrogen cryogen used to cool the instruments W3 and W4. Next, we limited our MIR study on the current sample to the W1 and W2 bands only due to the availability of the data for the W3 and W4 bands between 2010 and 2011, with many photometric points marked as ‘null’. Furthermore, both databases were screened to exclude bad photometric measurements, adopting the criteria (i) the reduced value of for the single exposure profile fit must be below 5 for both the W1 and W2 bands (that is, w1rchi2 5 and w2rchi2 5). (ii) The number of points spread function (PSF) components used in the profile fit for a source should be fewer than 3 (nb 3). (iii) Single-exposure images must have the highest quality (qifact = 1), be free from known artifacts (ccflags = ‘0000’), and not involve active d-eblending (na = 0). (iv) The number of photometric data points available in the W1 and W2 bands to be 5. (v) Also removed outliers by applying a 3 clipping on the whole data set.
@@ 3 Methodology for analysis
Before proceeding with the analysis further, we first corrected the observed data of each source for the galactic interstellar reddening and absorption foreground, using the extinction correction (AV) values provided by NASA’s NED, following the method of (Schlafly & Finkbeiner, 2011). The extinction-corrected r-band light curves for a gNLS1, ngNLS1, and gBLS1 from the current sample are shown in Fig. 2.
@@ 3.1 Optical flux variability
To confirm the presence/absence of variability in the final samples of gNLS1s, ngNLS1s, and gBLS1s (see Table 1) using the ZTF g, r, and i band light curves, we have used the standard -test (see, Goyal et al., 2012). A detailed explanation for this test is demonstrated in our previous papers (Ojha et al., 2022; Ojha et al., 2024, and references therein).
In summary, the following Goyal et al. (2012), -test can be written as
@@
where, is the mean square error and is given as follows:
is the sample variance for the flux measurements , . The associated error with can be computed following the given equation.
For each gNLS1, ngNLS1, and gBLS1, we have tabulated the duration of the light curve, the number of data points in the light curve, the variability statistics, based on the test , and in Table 2.
@@ 3.1.1 Duty cycle of variability
AGNs do not necessarily exhibit flux variations during each epoch of their monitoring. Hence, it is more suitable to evaluate the duty cycle (DC) as the ratio of the time intervals when variability is observed to the total observing time rather than simply considering the fraction of variable objects. Following the definition proposed by Romero et al. (1999) and further utilized by Stalin et al. (2004) for intra-night variability, we computed the DC for the current sample of Seyfert galaxies using the following expression
@@ 3.2 Mid-infrared flux variability
where is the magnitude observed at the point and is the weighted average. The can be expressed as:
where is calculated including the systematic error () with the individual error () as follows: